OpenStudy (anonymous):

Multiply. \frac{3}{5} x \frac{2}{9} A. \frac{2}{45} B. \frac{5}{14} C. \frac{2}{15} D. \frac{2}{3}

OpenStudy (anonymous):

@PokemonFan40

OpenStudy (pokemonfan40):

yeah?

OpenStudy (anonymous):

Multiply. \frac{3}{5} x \frac{2}{9} A. \frac{2}{45} B. \frac{5}{14} C. \frac{2}{15} D. \frac{2}{3}

OpenStudy (anonymous):

@just_one_last_goodbye

OpenStudy (anonymous):

that's not helpful..?

OpenStudy (just_one_last_goodbye):

Need help?

OpenStudy (pokemonfan40):

go there to see if it can help you

OpenStudy (anonymous):

I do need help D:

OpenStudy (anonymous):

site is not helpful though

OpenStudy (lolacole12):

(35x)(2)9 =215x(Decimal: 0.133333x)

OpenStudy (anonymous):

what does that mean?>

OpenStudy (just_one_last_goodbye):

@AnimeWover your answer for #1 ^_^ http://i.imgur.com/t7ZTiGa.png

OpenStudy (anonymous):

it's 3 - 5 not 35

OpenStudy (anonymous):

ty. where did u find that? 0-0

OpenStudy (just_one_last_goodbye):

I have a special program I designed ^_^ explain it faster

OpenStudy (anonymous):

XD

OpenStudy (just_one_last_goodbye):

Please medal me if you found it helpful

OpenStudy (anonymous):

Could you help me more?

OpenStudy (just_one_last_goodbye):

sure make a new post ^_^